OpenStudy (anonymous):

A wooden artifact from an ancient tomb contains 70% of the carbon-14 that is present in living trees. How long ago was the artifact made? (The half-life of carbon-14 is 5730 years. Round your answer to the nearest whole number.)

OpenStudy (anonymous):

You know that every 5730 years, half of the carbon-14 decays. So the equation per year would be (1/2)^(x/5730). Now, to find 70 percent of the C-14 that has decayed, you would assign 100 as the original amount of carbon and 70 as the current amount. So...

OpenStudy (anonymous):

70=100*(1/2)^(x/5730)

OpenStudy (anonymous):

Now just solve for x

OpenStudy (anonymous):

0.7=(1/2)^(x/5730)

OpenStudy (anonymous):

log(base1/2)(0.7)=x/5730

OpenStudy (anonymous):

x/5730 =log (base 10)(0.7)/ log(base 10) (0.5)=0.5146 so, x = 0.5146*5730=2948.5 or 2948 years